OpenStudy (anonymous):

Simplify -2[9 - (x + 7)].

OpenStudy (anonymous):

2x-4

OpenStudy (decentnabeel):

\[-2\left(-\left(x+7\right)+9\right)=-2\left(-x+2\right)\]

OpenStudy (anonymous):

?

OpenStudy (decentnabeel):

first solve the parenthesis() the slove the square bracket[]

OpenStudy (anonymous):

oh

OpenStudy (usukidoll):

the negative isn't being distributed correctly on the second to the last line -2[2-x]

OpenStudy (usukidoll):

while -2 x 2 = -4 a negative times a negative is positive (-2)(-x)

OpenStudy (decentnabeel):

-2[9 - (x + 7)] -2[9-x-7] -2[2-x] -4+2x

OpenStudy (usukidoll):

much better ^

OpenStudy (decentnabeel):

thanks @UsukiDoll